# When LLMs Play the Telephone Game: Cumulative Changes and Attractors in Iterated Cultural Transmissions

### v1

@@ 4 Results
For each of the 5 models, 3 tasks, and 20 initial texts, we ran 5 transmission chains with 50 transmission steps. We provide some examples of generated texts in Appendix Section 1, and complete data on the companion website 888https://sites.google.com/view/telephone-game-llm. By extracting the properties of generated texts at each generation of each chain, we can study the evolution of these properties through generations, measure how they are affected by interactions beyond single-turn effects, as well as detect and characterize theoretical attractors. By comparing the semantic similarities of texts produced by different chains, we can also evaluate whether sets of chains tend to converge or diverge.
@@ 4.1 Qualitative analysis of property evolutions over generations
We first hypothesized that iterated transmissions would affect the properties of generated texts beyond the first transmission. Across many instances, we indeed found that text properties keep evolving after the first generation. In Figure 3, we show the evolution of text difficulty, positivity, length and toxicity for one of the 20 initial texts, for all models and all tasks. Evolution of these properties for each of the 20 initial texts can be found in Appendix Section B. This specific example already allows to notice important differences in dynamics between models, tasks, and properties. Indeed, we observe that while toxicity converges to values close to zero for all models and tasks, this happens at a slower pace for GPT3.5 in the Rephrase task, and for Llama3-70B in the Take Inspiration task. For positivity, we observe that on the Take Inspiration and Continue tasks, GPT3.5 and Mixtral-8x7B converge to high positivities almost instantly, while evolution is more gradual for other models. The dynamics of difficulty appears to be highly influenced by task and models, as we observe cumulative changes for Llama3-8B and Llama3-70B in the Take Inspiration and Continue tasks, but not so much for other tasks and models. Length appears to exhibit little to no cumulative dynamics in this example. Interestingly, there seems to be a discontinuity of high magnitude for Mistral-7B toward the end of the chain in the Continue task. Qualitative observation of the texts revealed that this appears to be an example of collapsing behavior, which we discuss more extensively in Appendix Section B.
Overall, these results suggest that the evolution of text properties across repeated transmission is highly sensitive to both agent models and tasks.
Although we here focused on a specific example, looking at the evolution over generations of the distribution of each properties gives a higher-level idea of the dynamics. In Figure 4, we show the evolution of property distributions over generation for each model and task. This reveals important difference depending on the analyzed property, the task and the model. For instance, we observe that toxicity (Figure 4.a ) converges very quickly to a very narrow peak centered around 0. This is very different from the evolution of positivity (Figure 4.b), for which the initial distribution appears to be quite preserved for the Rephrase task (Figure 4.b, first row), while less constrained tasks such as Take inspiration (second row) and Continue (third row) lead to more visible changes. Interestingly, we observe that for Llama3-8B (blue) and Llama3-70B (pink), the distribution of positivity values converges to a bimodal distribution, while distributions are unimodal for other models. In some cases, we also observe that different models lead the distributions to be shifted in opposite directions. For instance when looking at the evolution of texts length (Figure 4.d), using GPT3.5 (green) or Llama3-8B (blue) leads text to become on average shorter, while using Mixtral-8x7B shifts the distribution towards greater length values.
@@ 4.2 To what extent do multi-turn transmissions affect the evolution of properties?
Qualitative analyses from the previous section appear to suggest that multi-turn transmissions lead texts to acquire different properties compared to single-turn settings. To quantitatively evaluate this observation, we use Kolmogorov–Smirnov (KS) tests [41] to estimate the compare property distributions after a single interaction and after multiple interactions. In Figure 5, we report for each model, task and property the p-value of the KS test for the null hypothesis : "The text properties at generation are sampled from the same distribution as the text properties after generation 1". Across most instances, we observe that can be confidently rejected, indicating that property distributions become significantly different after multi-turn transmissions compare to single-turn transmissions. We observe that this is more often the case for less constrained tasks (Take Inspiration and Continue, second and third columns) than for more constrained task (Rephrase, first column). This finding confirms that studying single-turn interactions is in general not sufficient for analyzing the property of interacting LLMs outputs. This warrants a more detailed account of the cultural dynamics across iterated interactions among LLMs.
@@ 4.3 What influences the presence, strength, and position of attractors?
Visual inspection of the evolution of text properties as presented on Figure 4 indicate that multi-turn transmissions lead distributions to become skewed toward certain values, which suggests the presence of attractors. The task assigned to a chain and the model type populating it appear to influence the position of those attractors, as well as their strength (i.e. how quickly do shifts in distributions happen). To have quantitative measures of attractors strengths and positions, we use the method described in 3.4 and Figure 2. By fitting a linear regression predicting the value of a given property at the end of the chain as a function of the property in the initial text, we can represent for each property, model and task the position and strength of the attractors, provided they exist. Figure 6 presents the estimated strengths and positions of attractors, and fitted linear regressions are provided as supplementary material in Figure 8.
@@
For all combinations of property, task and model, we found that the recurrent relationship defined by the fitted linear regression converges. This means that all conditions admit a theoretical attractor as defined in Section 3.4. As for the strength and position of these theoretical attractors, Figure 6 already allows to notice some tendencies. For instance, it seems that less constrained tasks (e.g. Continue) lead to stronger attractors than more constrained tasks (e.g. Rephrase). To better disentangle the respective contributions of model type, task and property on attractors, we fitted bayesian models predicting attractor Strength as a function of Task, Model and Property, and predicting attractor Position as a function of Task and Model, for each of the four considered properties. Details about statistical analyses are provided in Appendix section B.
@@
We find a strong effect of Task on attractor strength: Continue leads to significantly stronger attraction than Take Inspiration, itself leading to significantly stronger attraction than Rephrase. This confirms our observation that less constrained tasks lead to stronger attraction than more constrained tasks. Different properties are also find to display different sensitivity to attraction effects. We detect that toxicity possesses significantly stronger attractors than positivity, difficulty and length. As for the effect of model, we observe significantly weaker attraction for Llama3-70b compared to GPT3.5, Llama3-8B and Mixtral-8x7b.
@@
As for the position of the attractors, we observe the attractor for toxicity is significantly higher for Llama3-8b than for GPT3.5 and Mixtral-8x7b, and significantly higher for Llama3-70b than for GPT3.5, Mistral-7b and Mixtral-8x7b. It is also higher for Continue than for Rephrase. For positivity, we found that the position of the attractor was significantly lower for Llama3-8b than for GPT3.5, Mistral-7b and Mixtral-8x7b, and that the task Take inspiration and Continue both led to significantly higher positivity than the Rephrase task.
4.4 To what extent do different transmission chains converge on similar content?
Lastly, we investigate the extent to which iterated transmissions lead different chains to diverge or converge as determined by whether the between-chain similarity among generated texts increases or decreases after several generations. This can be measured by assessing the cosine similarity between final text embeddings versus the cosine similarity between initial text embeddings for each possible pair of chains, for each model and task (Figure 7).
As expected given the nature of the task, chains instructed with the Rephrase task maintained close similarity with the initial text (Figure 7). Out of the five model tested, Llama3-8B was more likely to maintain semantic similarity across generations for this prompt. For Take inspiration and Rephrase, there seems to be a tendency for chains to lead to a specific distance between final texts, as the initial distance between texts has little impact on the final distance. This means that over generations, chains that started with very similar texts diverge while chains that started with very different texts converge. The position of this attractor appears to be influenced by the model, as for example Llama3-8b displays much more convergence than the two other models for the Take inspiration task.
@@ 5 Discussion
By comparing the properties of input texts to those of texts produced by transmission chains spanning several generations of LLMs, we identified property-specific patterns in the convergence of LLM dynamics toward attractor states. Although the existence of fixed points identified with our method does not prove the existence of an attractor, it allows to perform quantitative evaluations on the dynamics of text evolution across models, tasks and properties. Using this method, we found high convergence rates for some text properties (toxicity and positivity), independent of the chain’s model and task whereas the evolution of other text characteristics (difficulty and length) was influenced by the task and the model. Differences in dynamics among properties could be a consequence of fine-tuning through reinforcement learning with human feedback, a commonplace practice in LLM training which may target some properties more than other (e.g., specifically avoiding toxic content without addressing its difficulty), creating strong attractors. It is also interesting to notice that convergence rates were, on average, higher for more open-ended tasks (Take inspiration and Continue) than a more constrained one (Rephrase), suggesting the study of cultural evolution in LLM transmission chains might be particularly relevant to situations in which LLMs are used to simulate artificial societies [20, 51, 52, 77, 34, 71, 17], where they are often granted a relatively high freedom in order to witness emergent behaviors.
We also introduced several evaluation metrics for analyzing cultural dynamics, in particular defining a task- and metric-independent notion of theoretical attractor. Although these methods were developed to study transmission chains in LLMs, similar tools may be applied to studies of human cultural evolution, allowing for inferences across tasks and cultural domains, moving beyond domain-specific results.
@@
As the study of the cultural dynamics of generative agents is an emerging research area, our setting involved several simplifications. While we focused on linear transmission chains, real-world interactions typically involve networks of senders and receivers. Human studies have shown that network size [28, 60, 4, 6, 21] and structure [58, 38, 22, 19, 23] influence cultural evolution. Following some initial endeavors [50, 53], future work may assess similar effects in machine networks. To investigate model- and task-specific biases, we studied transmission chains in homogeneous settings, where agents belonged to the same model type and received the same instructions, but future studies may address cultural dynamics in heterogeneous populations of LLMs prompted with various instructions. While we focused on LLM and, therefore, text outputs, similar studies may be run to address the properties of various generative tools (e.g., for image generation). In the future, researchers may also address hybrid networks in which humans and LLMs interact — a scenario that is becoming increasingly relevant as generative tools become more widespread and which may, in turn, shape the future of human cultural evolution. [10].
Acknowledgements
This research was partially funded by the French National Research Agency (ANR, project ECOCURL, Grant ANR-20-CE23-0006). This work benefited from access to the Jean Zay (Idris) supercomputer associated with the Genci grant A0151011996. We also thank Chris Foulon, Marcela Ovando-Tellez and Joan Dussauld who participated in the hackathon Hack1Robo during which this project originated.
@@ Appendix A Additional details on the methods
Data analyses revealed that, on the Continue task, when using Mistral-7B, agents of the chains would sometimes start outputting very long text by filling them with “#some_keyword”. As this behavior created a few outliers, we thought it would be better to filter-out those “#some_keyword” when performing the main analyses. This behavior is nevertheless an interesting result, reminiscent of the collapsing dynamics found when training LLMs on their own output [65]. We therefore discuss it separately in Appendix B.
@@
Positivity. We employ the SentimentIntensityAnalyzer tool from the NLTK library to assess the positivity of generated texts. The tool is based on the Valence Aware Dictionary and sEntiment Reasoner (VADER) method [35], which is a lexicon and rule-based sentiment analysis tool specifically designed for social media data. It uses a combination of lexical features, such as words and their semantic orientation, to determine the overall sentiment of a text. In the VADER method, every word in the vocabulary is rated with respect to its positive or negative sentiment and the intensity of that sentiment. The SentimentIntensityAnalyzer uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
@@ Appendix B Additional figures and analyses
We performed statistical analyses using the Python package pymc [75] to fit Bayesian models.
@@
To determine the significance of the difference between estimated parameters, we computed the 95% credibility intervals of the difference by sampling from the posteriors. In Figures 9 and 10, we provide the matrices indicating the mean of this difference between estimated parameters. Stars indicate that the credibility interval does not contain 0.
@@
Figure 11 shows the average length of text generated with the Mistral-7B model chain on the Continue task for five different seeds of the same story. We can observe several discontinuities in terms of the generated text length, i.e. at some iterations the length drastically increases or decreases. It is interesting to note that when the length decreases, it returns to the original value as before the first discontinuity. This suggests the existence of an attractor regarding this specific length. To better understand the cause of these discontinuities, Figure 2 shows examples of stories generated before and after those discontinuities (for the seed number three in figure 11). We can see that at generation 14 the model abruptly starts to generate many hashtags. It generates 283 hashtags, compared to 12 in the previous generation. At generation 45, we can see that the overall quality of the text decreased into generating solely hashtags and brief descriptions. This reduction in text quality is reminiscent of collapsing dynamics observed in iterative chains of LLMs, where each model was trained on the output of a previous one [65].
@@
The method introduced in Section 3.4 gives the position and strength of a theoretical attractor (or theoretical fixed point). In order to validate our method, we verified that this theoretical prediction matches the actual data. To do so, we used the first 10 generations of each simulated chain to predict the strength and position of attractors for each task, model and property. We then compared this prediction with the actual properties of texts obtained after 50 generations. As shown in Figure 12, transmission chains shifts the initial distribution of values in the direction of the predicted attractor. Moreover, the variance of the final distribution appears to reflect the predicted strength of the attractor. These results confirmed that the method we introduce is indeed suited for estimating the strength of position of attractors.

### v2

@@ 4 Results
For each of the 5 models, 3 tasks, and 20 initial texts, we ran 5 transmission chains with 50 transmission steps. We provide some examples of generated texts in Appendix 1, and complete data on the companion website 888https://sites.google.com/view/llms-play-telephone . By extracting the properties of generated texts at each generation of each chain, we can study the evolution of these properties through generations, measure how they are affected by interactions beyond single-turn effects, as well as detect and characterize theoretical attractors.
@@ 4.1 Qualitative analysis of property evolutions over generations
In Figure 2, we show the evolution of property distributions over generation for each model and task. This reveals important difference depending on the analyzed property, the task and the model. For instance, we observe that toxicity converges very quickly to a very narrow peak centered around 0. This is very different from the evolution of positivity, for which the initial distribution appears to be quite preserved for the Rephrase task (Figure 2, first row), while less constrained tasks such as Take inspiration (second row) and Continue (third row) lead to more visible changes. Interestingly, we observe that for Llama3-8B (blue) and Llama3-70B (pink), the distribution of positivity values converges to a bimodal distribution, while distributions are unimodal for other models. In some cases, we also observe that different models lead the distributions to be shifted in opposite directions. For instance when looking at the evolution of texts length, using GPT3.5 (green) or Llama3-8B (blue) leads text to become on average shorter, while using Mixtral-8x7B shifts the distribution towards greater lengths.
@@ 4.2 To what extent do multi-turn transmissions affect the evolution of properties?
Qualitative analyses from the previous section appear to suggest that multi-turn transmissions lead texts to acquire different properties compared to single-turn settings. To quantitatively evaluate this observation, we use Kolmogorov–Smirnov (KS) tests ks_test to estimate the compare property distributions after a single interaction and after multiple interactions. In Figure 3, we report for each model, task and property the p-value of the KS test for the null hypothesis : "The text properties at generation are sampled from the same distribution as the text properties after generation 1". Across most instances, we observe that can be confidently rejected, indicating that property distributions become significantly different after multi-turn transmissions compare to single-turn transmissions. We observe that this is more often the case for less constrained tasks (Take Inspiration and Continue, second and third columns) than for more constrained task (Rephrase, first column). This finding confirms that studying single-turn interactions is in general not sufficient for analyzing the properties of interacting LLMs’ outputs. This warrants a more detailed account of the cultural dynamics across iterated interactions among LLMs.
@@ 4.3 What influences the presence, strength, and position of attractors?
Visual inspection of the evolution of text properties as presented on Figure 2 indicate that multi-turn transmissions lead distributions to become skewed toward certain values, which suggests the presence of attractors. The task assigned to a chain and the model type populating it appear to influence the position of those attractors, as well as their strength (i.e. how quickly do shifts in distributions happen). To have quantitative measures of attractors strengths and positions, we use the method described in 3.4 and Figure S1. By fitting a linear regression predicting the value of a given property at the end of the chain as a function of the property in the initial text, we can represent for each property, model and task the position and strength of the attractors, provided they exist. Figure 4 presents the estimated strengths and positions of attractors, and fitted linear regressions are provided as supplementary material in Figure S2.
@@
For all combinations of property, task and model, we found that the recurrent relationship defined by the fitted linear regression converges. This means that all conditions admit a theoretical attractor as defined in Section 3.4. As for the strength and position of these theoretical attractors, Figure 4 already allows to notice some tendencies. For instance, it seems that less constrained tasks (e.g. Continue) lead to stronger attractors than more constrained tasks (e.g. Rephrase). To better disentangle the respective contributions of model type, task and property on attractors, we fitted Bayesian models predicting attractor Strength as a function of Task, Model and Property, and predicting attractor Position as a function of Task and Model, for each of the four considered properties. We report 95% Credible Intervals (CI), and provide details about statistical analyses in Appendix B.
@@
We find a strong effect of Task on attractor strength: Continue leads to significantly stronger attraction than Take Inspiration (95% CI: [0.0511;0.254]) , itself leading to significantly stronger attraction than Rephrase (95% CI: [0.0326; 0.2377]). This confirms our observation that less constrained tasks lead to stronger attraction than more constrained tasks. Different properties are also found to display different sensitivity to attraction effects. We detect that toxicity possesses significantly stronger attractors than positivity (95% CI: [0.1158; 0.3515]), difficulty (95% CI: [0.1072; 0.3417]) and length (95% CI: [0.173; 0.407]). As for the effect of model, we observe significantly weaker attraction for Llama3-70b compared to GPT3.5 (95% CI: [-0.3215; -0.059]) , Llama3-8B (95% CI: [-0.2954; -0.0326]) and Mixtral-8x7b (95% CI: [-0.3163; -0.0541]) .
@@
We next analyze the position of the attractors. We observe that the attractor for toxicity is significantly higher (i.e the attractor state has higher toxicity) for Llama3-8b than for GPT3.5 and Mixtral-8x7b, and significantly higher for Llama3-70b than for GPT3.5, Mistral-7b and Mixtral-8x7b. It is also higher for Continue than for Rephrase. For positivity, we found that the position of the attractor was significantly lower for Llama3-8b than for GPT3.5, Mistral-7b and Mixtral-8x7b, and that the task Take inspiration and Continue both led to significantly higher positivity than the Rephrase task.
Effect of temperature
Temperature is a hyper-parameter that controls the stochasticity in LLMs outputs. Our experiments indicate that increasing the temperature appears to lead to stronger attractors only in constrained tasks (Rephrase and Take inspiration), but not in the more open- ended task Continue (Figure 5a). One interpretation might be that increasing temperature relaxes constraints on the content that can be produced, thus leading to stronger attractors. This effect would therefore be more significant for tasks that are quite constrained.
Effect of fine-tuning
We observe that fine-tuning can shift the attractors’ positions (Figure 5b): the attractor for length is lower, and the attractor for difficulty higher, for Instruct models compared to Base models. For toxicity and positivity, we do not observe very significant shifts. This may indicate that Base models were already quite aligned with human preferences in terms of toxicity and positivity even before fine-tuning, possibly due to techniques such as data curation. Length and Difficulty may have been less targeted by such techniques, and fine-tuning may therefore have shifted the attractors’ positions towards human preferences. As for the strength of attractors, we observe that fine-tuning seems to increase attraction strength for toxicity, but to reduce it for difficulty. This may suggest that for properties on which humans have strong, uniform preferences (such as toxicity), fine-tuning leads to stronger attractors, while for properties on which humans have weaker and more heterogeneous preferences, it mitigates the strength of attractors that came from the training data. This analysis provides insight about how attractors are formed, and reveals that both the choice of training data and fine-tuning processes may impact LLMs in ways that only become significant in the case of multi-turn behaviors.
@@ 5 Discussion
Those qualitative observations were confirmed by statistical tests, where we found that multi-turn interactions lead the distributions of text properties to become significantly different distributions obtained after single-turn interactions.
By comparing the properties of input texts to those of texts produced by transmission chains spanning several generations of LLMs, we identified property-specific patterns in the convergence of LLM dynamics toward attractor states. Using this method, we found that some properties (e.g. toxicity) display quicker convergence rates toward attractors, and that the position of these attractors varies between models. Moreover, our results suggest that more open-ended tasks (e.g. Continue) lead to quicker convergence toward attractors than more constrained tasks (e.g. Rephrase). Studying attraction dynamics in systems of LLMs may therefore be particularly relevant to situations in which LLMs are used to simulate artificial societies de_zarza_emergent_2023; park_generative_2023; park_social_2022; xiao_simulating_2023; hua_war_2023; vezhnevets_generative_2023; chuang2023simulating, where they are often granted a relatively high freedom in order to witness emergent behaviors. Evaluating the effect of temperature on attractors supports this view that allowing more room for variation leads to stronger attractors, as we found that increasing temperature increases attraction strength only for quite constrained tasks.
To understand the role of fine-tuning in shaping the strength and position of these attractors, we compared Base models with their Instruct versions. This revealed that fine-tuning can affect both attractors strength (e.g. for toxicity) and position (e.g. for difficulty and length).
The main practical consequence of our results is that they question the current ways in which LLMs are fine-tuned and evaluated. Indeed, current fine-tuning procedures, such as RLHF, rely on a single interaction to rate the output of the LLMs. Similarly, the vast majority of benchmarks and leaderboards are based on the capacities of LLMs in single-turn interactions. As multi-step interaction is a major use-case of LLMs (e.g. when interacting iteratively with a chatbot, or when LLMs are used in multi-agent simulations), there is a contradiction between how LLMs are used and how LLMs are evaluated and trained. Our results indicate that this contradiction cannot be ignored: for example, on Figure S24 (toxicity - take inspiration) we see that at first Mistral appears less toxic at the first iteration, but after 9 generations Llama3-8B ends up being much less toxic. One may therefore choose Mistral-7B based on single-step evaluation, even though Llama3-8B is the optimal choice for many applications.
More generally, our results reveal that knowing the output of LLMs after a single interaction is insufficient to predict their behavior in long-term social interactions. Therefore, this suggests that existing training and evaluation methods are not suited to align the behavior of LLMs in multi-step settings. The concepts of attractor position and strength may offer a way to take into account the consequence of multi-turn interactions when comparing and evaluating LLMs.
@@
As the study of the cultural dynamics of generative agents is an emerging research area, our setting involved several simplifications. While we focused on linear transmission chains, real-world interactions typically involve networks of senders and receivers. Human studies have shown that network size fay_increasing_2019; richerson_group_2013; andersson_group_2014; baldini_revisiting_2015; derex_experimental_2013 and structure raviv_role_2020; kirby_cumulative_2022; derex_partial_2016; de_pablo_understanding_2022; derex_cumulative_2020 influence cultural evolution. Following some initial endeavors nisioti_social_2022; perez_cultural_2024, future work may assess similar effects in machine networks. To investigate model- and task-specific biases, we studied transmission chains in homogeneous settings, where agents belonged to the same model type and received the same instructions, but future studies may address cultural dynamics in heterogeneous populations of LLMs prompted with various instructions. While we focused on LLM and, therefore, text outputs, similar studies may be run to address the properties of various generative tools (e.g., for image generation). In the future, researchers may also address hybrid networks in which humans and LLMs interact — a scenario that is becoming increasingly relevant as generative tools become more widespread and which may, in turn, shape the future of human cultural evolution. machine_culture.
@@ Appendix A Additional details on the methods
Data analyses revealed that, on the Continue task, when using Mistral-7B, agents of the chains would sometimes start outputting very long text by filling them with “#some_keyword”. As this behavior created a few outliers, we thought it would be better to filter-out those “#some_keyword” when performing the main analyses. This behavior is nevertheless an interesting result, reminiscent of the collapsing dynamics found when training LLMs on their own output [shumailov2023curse]. We therefore discuss it separately in Appendix B.
@@
Environmental footprint
Taking heat recovery in account, running experiments consumed about 777kWh, which represents about 15.54 kg of CO2.
@@
Positivity. We employ the SentimentIntensityAnalyzer tool from the NLTK library to assess the positivity of generated texts. The tool is based on the Valence Aware Dictionary and sEntiment Reasoner (VADER) method [hutto2014vader], which is a lexicon and rule-based sentiment analysis tool specifically designed for social media data. It uses a combination of lexical features, such as words and their semantic orientation, to determine the overall sentiment of a text. In the VADER method, every word in the vocabulary is rated with respect to its positive or negative sentiment and the intensity of that sentiment. The SentimentIntensityAnalyzer uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
@@ Appendix B Additional figures and analyses
We performed statistical analyses using the Python package pymc [wiecki_pymc-devspymc_2024] to fit Bayesian models.
@@
To determine the significance of the difference between estimated parameters, we computed the 95% credibility intervals of the difference by sampling from the posteriors. In Figures S3 and S4, we provide the matrices indicating the mean of this difference between estimated parameters. Stars indicate that the credibility interval does not contain 0.
@@
Figure S5 shows the average length of text generated with the Mistral-7B model chain on the Continue task for five different seeds of the same story. We can observe several discontinuities in terms of the generated text length, i.e. at some iterations the length drastically increases or decreases. It is interesting to note that when the length decreases, it returns to the original value as before the first discontinuity. This suggests the existence of an attractor regarding this specific length. To better understand the cause of these discontinuities, Figure 2 shows examples of stories generated before and after those discontinuities (for the seed number three in figure S5). We can see that at generation 14 the model abruptly starts to generate many hashtags. It generates 283 hashtags, compared to 12 in the previous generation. At generation 45, we can see that the overall quality of the text decreased into generating solely hashtags and brief descriptions. This reduction in text quality is reminiscent of collapsing dynamics observed in iterative chains of LLMs, where each model was trained on the output of a previous one [shumailov2023curse].
@@
The method introduced in Section 3.4 gives the position and strength of a theoretical attractor (or theoretical fixed point). In order to validate our method, we verified that this theoretical prediction matches the actual data. To do so, we used the first 10 generations of each simulated chain to predict the strength and position of attractors for each task, model and property. We then compared this prediction with the actual properties of texts obtained after 50 generations. As shown in Figure S6, transmission chains shifts the initial distribution of values in the direction of the predicted attractor. Moreover, the variance of the final distribution appears to reflect the predicted strength of the attractor. These results confirmed that the method we introduce is indeed suited for estimating the strength of position of attractors.